OpenStudy (anonymous):

A block of certain radioactive substance having mass of 140g is observed. This radioactive substance is known to decay at a rate proportional to the amount present. After 50 hours, its mass reduces to 120g, find 1. an expression for the mass of the substance at any time. 2. The time-lapse before the block decays to one quarter of its original mass.

OpenStudy (anonymous):

\[120\div140=\frac{6}{7}\] so you can use \[A(t)=140\times (\frac{6}{7})^{\frac{t}{50}}\] if you like

OpenStudy (anonymous):

you are probably supposed to use \[A(t)=140\times e^{kt}\] and find k and all that work, but it is unnecessary

OpenStudy (anonymous):

if you want 1/4 =.25 of the original amount set \[.25=(\frac{6}{7})^{\frac{t}{50}}\] and solve for t via \[\ln(.25)=\frac{t}{50}\ln(\frac{6}{7})\] and do \[t=\frac{50\ln(.25)}{\ln(\frac{6}{7})}\]